OpenStudy (anonymous):

using the function f(x)=2x and g(x)=xsquare +1 to find the value of g(3)+f (4)

OpenStudy (anonymous):

f(4)=2.4=8 and g(3)= (3)^2+1=10. g(3)+f(4)=10+8=18